40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 11 of 56  Not logged in ELOG logo
ID Date Author Type Category Subject
  2305   Wed Mar 13 12:44:41 2019 awade, anchalDailyProgressWOPOPol launch into PM fiber 1064 nm

[awade, anchal ]

After a bit of reading I've realized that the standard use of these PM fibers is to launch along the slow axis (see for example Thorlabs and OzOptics info on fiber beam splitters).  It should be much of the sameness for patch cables, but polarization sensitive elements like beam splitters are mostly tested and specified for slow axis launch unless they are custom made to order. 

We are switching the polarization alignment to slow axis in the 1064 nm and 532 nm fiber coupling.  Anchal is re-optimizing​ the 1064 nm launch to get the PM fiber extinction ratio back to a good place.  We've also changed input launch to use a laser line PBS mounted in a rotation mount for clean linear polarization.  With the optimized setup the for the 1064 nm fiber path the output polarization signal goes from 3700 mV to 39.3 mV which is an extinction​ ratio of -19.7 dB.

Here the max theoretical extinction​ ratio is 

ER = -10 \log_{10}[\tan^2(\theta)]

which would place our goodness of alignment to with 0.61 deg.

Updated 1064 nm launch. Uses rotation mounted PBS for guaranteed linear
polarization, half wave plate is to maximize power.

 

 

Quote:

I've replaced the SM fiber in the 1064 nm launch with a PM fiber (P3-1064PM-FC-5). I also moved the fiber collimator (F240APC-1064) back 2.54 cm back to give more space for a PBS cube (to check linearly of the light).  

For the 1064 nm launch it seemed to be a lot harder to find the initial alignment of the collimator using the alignment of the back propagated 650 nm fiber laser source. Here I aligned a pair of irises in the forward propagating direction and then back propagated through the PM fiber using 650 nm to get the initial​ pointing of collimator. I don't know why this is so much harder than the 532 nm case.  I suspect one of the steering mirrors is not really reflecting off the front dielectric surface.  In the end I did a bunch of systematic walking of the fiber launch mount and eventually fount the alignment.  

From 4.44 mW of input light I get 2.74 mW of light out the other end of the fiber.  This is an efficiency of 62 % which is more than enough for my needs.  I expect the HD will only need 1 mW (2 mW max), so this is fine. Getting this in coupling higher will require a bit of lens walking, not really worth it at this stage.

I had already carefully aligned the collimator orientation to put the fast axis on aligned to p-pol (wrt the table), by eye.  It seems like the launch pretty much hit the correct launch polarization on the first go.  I see little variation in the polarization when I pulse the heat on the fiber.  This is now good to go for optimizing the homodyne visible and polarization overlap output from the SQZ.

 

  2304   Tue Mar 12 16:58:50 2019 awadeMiscWOPOTodo list W11

These are things to get do this week on WOPO experiment:

☑️ Reinstall 50:50 fiber splitter into homodyne setup and go back to fiber launch of both ends of the HD directly onto photodetectors (rather than free space)

☐ Check visibility of HD by launching 1064 nm into both arms of HD using splitter and extra length on one arm, ramp laser frequency to get fringes and look at Vpp on each diode seperatly

☑️ Optimize subtraction of HD for max Common Mode Rejection (CMR) of LO amplitude noise.  Inject 3.21 kHz line into laser BNC port and minimize this peak on the subtracted output of the HD.

☐ Check 1064 nm -> 532 nm conversion in WOPO device to establish polarization basis for correct pol alignment into fiber (change 532 nm launch polarization if necessary), should be along the fast axis but for some reason this isn't in the datasheet explictly

☐ Double check how much power change there is from 2 pi modulation depth from PZT mounted 1064 nm mirror.  We don't want to over actuate on this element as it slightly misaignes into the fiber launch and causes some change in power, this is ok over a small range as the CMR of the homodyne will reject this.  We just want to be sure that this isn't the dominant effect that we are seeing at the output once we thing we should be seeing SQZ.

 

☑️Tick mark when done.

  2303   Tue Mar 12 16:35:43 2019 awadeDailyProgressWOPOPol launch into PM fiber 1064 nm

I've replaced the SM fiber in the 1064 nm launch with a PM fiber (P3-1064PM-FC-5). I also moved the fiber collimator (F240APC-1064) back 2.54 cm back to give more space for a PBS cube (to check linearly of the light).  

For the 1064 nm launch it seemed to be a lot harder to find the initial alignment of the collimator using the alignment of the back propagated 650 nm fiber laser source. Here I aligned a pair of irises in the forward propagating direction and then back propagated through the PM fiber using 650 nm to get the initial​ pointing of collimator. I don't know why this is so much harder than the 532 nm case.  I suspect one of the steering mirrors is not really reflecting off the front dielectric surface.  In the end I did a bunch of systematic walking of the fiber launch mount and eventually fount the alignment.  

From 4.44 mW of input light I get 2.74 mW of light out the other end of the fiber.  This is an efficiency of 62 % which is more than enough for my needs.  I expect the HD will only need 1 mW (2 mW max), so this is fine. Getting this in coupling higher will require a bit of lens walking, not really worth it at this stage.

I had already carefully aligned the collimator orientation to put the fast axis on aligned to p-pol (wrt the table), by eye.  It seems like the launch pretty much hit the correct launch polarization on the first go.  I see little variation in the polarization when I pulse the heat on the fiber.  This is now good to go for optimizing the homodyne visible and polarization overlap output from the SQZ.

  2302   Mon Mar 11 08:49:11 2019 Aidan, ChubComputingCymacsBuilding FB4 Cymacs

We started migrating equipment for the FB4 Cymacs to the QIL on Friday. See attached list and images.

Attachment 1: IMG_8851.jpg
IMG_8851.jpg
Attachment 2: IMG_8852.jpg
IMG_8852.jpg
  2301   Fri Mar 8 17:05:52 2019 AnjaliUpdate Frequency stabilization of 2 micron source

 

The schematic of the homodyne configuration is shown below.

Following is the component list

Item Quantity Availability Part number Remarks
Laser diode 1 Yes EP 2004-0-DM-B06-FA  
Isolator 1 Yes    
Coupler(50/50) 5 Yes TW2000R5A2A Fiber type : SM2000
Delay fiber   Yes SM2000 Need to check the exact length of the fiber
Photodetector 4 Yes DET10D  
SR560 3 Yes    

 

Attachment 1: Homodyne_setup_2micron.png
Homodyne_setup_2micron.png
  2300   Thu Feb 28 16:00:59 2019 awadeDailyProgressWOPOPol launch into PM fiber

I've set up a rotating PBS and half-wave plate to provide polarization adjustment into the 532 nm fiber without misalignment the spatial alignment.  Here I've used a PRM1 rotation mount with a SM1PM10 lens tube mount for beam cube prisms.  The lens tube mount is supposed to be for pre-mounted cubes but I've inserted some shims to hold it in place and it seems to work well like that.  It means I can get a nice clean linear polarization at all rotations.

After spatially aligning the input beam I stepped the rotation of the PBS (and accordingly the L/2 wave plate) and pulsed the temperature of the fiber using a heat gun.  After some walking I found that for the current fiber rotation (0 deg) the linear polarization was aligned with the fiber axis at 88 deg PBS rotation (here 0 deg PBS rotation is aligned for p-pol transmission, well almost). I made some adjustments to the alignment of the fiber collimator in the fiber launch, I aligned the slow axis key with the vertical so that the fast axis of the fiber is p-pol.

Keying on PM fibers

As a side note the keying of PM fiber patches is typically with the slow axis aligned with the key notch. The WOPO's PM fibers are keyed so that the alignment key is along the slow axis of the fiber (i.e. aligned with the stress rods). Figure below illustrates the configuration.

Replacing the 532 nm patch with fresh PM fiber

I was getting a large jitter in the power levels as measured at the output of the old SM and PM fibers (on the order of 10%).  These power fluctuations were not present on the input side.  I thought this was an alignment jitter or a polarization effect.  However, I was unable to minimize it by improving the input polarization at the launch.  When I tapped various mounts there didn't seem to be a corresponding correlation with output power jitter of the fiber.  When I checked the end of the PM fiber (P3-1064PM-FC-2), I saw that there was damage about the core (see pictured below).  It seems like maybe I had some kind of etalon effect from this burn mark and the launch.  After replacing the 532 nm PM fiber with a fresh one that arrived last week the power is much more stable and I was able to easily​ find the pol alignment going in. 

Damage to PM fiber end.  No amount of aggressive cleaning will remove the mark in the middle.
The fiber will need to be cut and a new connector spliced on.

 

Next job is to replace SM fiber for the 1064 nm delivery with PM fiber so there is a well defined polarization for launching into the homodyne detector.

 

 

Quote:

Alignment of the pumping 532 nm polarization into the WOPO is important to getting the correct phase matching condition.  For the periodically polled Lithium Niobate (LN) waveguide the phase matching is type-0: and pumping and fundamental wavelengths are in the same polarization.  The AdvR non-linear device is coupled with polarization maintaining fibers (Panda style), which are keyed at their FC/APC ends.  This means that with the correct launch polarization we should be correctly aligned with the proper crystal axis for degenerate down conversion (at the right chip temperature). 

Replacing Broadband PBS

Till now I was using non-pol maintaining patches to coupling into the WOPO fiber ends.  This should have been ok, but it is hard to figure out exactly which polarization is optimal so I switched to a pol-maintaining patch because it can be aligned separately and then the keyed connectors give you automatic alignment.  I had some issues trying to find the optimal polarization going into the fiber and I've now traced this back to the polarizing beam cubes.  I've been using Thorlabs PBS101 which is a 10x10x10 mm^3 beam cube that is supposed to be broad-band (420-680 nm).  When I checked the extinction ratio I saw Pmax=150 mW, Pmin=0.413 mW on transmission between extremes.  This is an extinction ratio of Tp:Ts = 393:1 which is much less than the spec of >1000:1.  Not sure what's going on here, the light going into the BS is coming directly from a Faraday isolator and a half-wave plate.  With some adjustment to the angle of the wave plate I can do a little better but it should be nicely linearly polarized to start with.

I've switched out the PSB101 for the laser line PBS12-1064 I remeasured extinction ratio (Pmax=150 mW, Pmin=27.6 µW) Tp:Ts = 5471:1 (better than the quoted 3000:1 spec).  This is good, at least now I know what is going on. I am also putting in an order for a 532 nm zero order quarter-wave plate, so that we can be absolutely sure we are launching in linear light always.  

Aligning light into pol-maintaining fiber

I previously thought I might be able to use the frequency modulation technique to align the light through the polarization maintaining fiber.  There is a birefringence in PM460-HP fiber of  3.5 x 10-4.  The phase between ordinary and extraordinary axes over the whole fiber length is

\Delta\phi = \frac{2 \pi \Delta n L}{c}f

Where L is fiber length, \Delta n is the birefringence and f is the laser frequency.  The idea is to launch linearly polarized light into the fiber and then at the readout place a polarizer rotated to be 90°: ramping frequency will produce an amplitude modulation on the dark fringe.  However, even with 1 GHz of frequency ramp this is only a 15 mrad effect for a 2 m fiber, its likely to be too small to see over other effects.  This is not enough to be able to fine align polarization.  

Instead I'll use the heat gun method.  I'll fire linearly polarized light into the fiber and measure the output with a crossed polarizer.  If the input polarization is correct there should be no power changes on the output as the fiber is thermally cycled. Its only two meters long so hopefully this effect is easy to see.

 

Attachment 2: PMFiber.pdf
PMFiber.pdf
  2299   Mon Feb 25 20:42:32 2019 KojiSummaryScatterometerSi Block stored in Cryo Lab

On the Friday cleaning, we vacated the east optical table. The Si scatterometer was disassembled and the Si block was moved and stored to the cryo lab.

Attachment 1: P_20190221_115501.jpg
P_20190221_115501.jpg
  2298   Fri Feb 22 14:05:09 2019 awadeDailyProgressWOPOMixing down homodyne: 5th order elliptic filter (100 kHz)

Making 5th order Elliptical Filter

I couldn't find any filters that would cut off above 100 kHz so I made my own using a Thorlabs EEAPCB1 generic filter PCB in a Thorlabs EEA14 enclosure. I used a 5th order elliptic design with a pass band up to 100 kHz and a stop band of 40 dB from 150 kHz.  To speed things up I used the Coilcraft Filter Designer software  (4.0.1, Windows) and chose closest standard values from parts we had in EE workshop kits. The Coilcraft designer is nice because it has the full physical model of the inductors built in.

The schematic is illustrated below:

5th order Elliptic Filter For use in homodyne readout electronics.  These are ideal values, actual values closest values were chosen from capacitor and inductor kits (listed in main text)

Actual values selected were closest available and I didn't try to do any mixing or matching to get fine tuned correct values. Values are as follows:

  • L2 and L4 27 µH => 1812CS-273XJL (27µH @ 2.5 MHz, tol 5%, Q_min 15 @ 2.5 MHz, SRF_min 11 MHz, DCR 11.04 Ω)
  • C1 20.47 nF => 12065C223KAT2A (0.022 µF, 50 V, 10%)
  • C2 14.38 nF => 12065153KAT2A (0.015 µF, 50 V, 10%)
  • C3 47.42 nF => 12061C473KAT2A (0.047 µF, 50 V, 10%)
  • C4 4.792 nF => 12065C472KAT2A (4700 pF, 50 V, 10%)
  • C5 27.3 nF => 12065C273KAT2A (0.027µF, 50 V, 10%)

The built filter is shown below.

The transfer function was taken from 10 Hz up to 5 MHz (IF BW of 10 Hz).  The Coilcraft Filter Designer software seems to export all the filter scattering parameters except S12 and S21, not the best. Instead I used LISO to model the filter's predicted response and this is plotted alongside measured TF.  Here I have scaled the filter model's response by 2 to match the impedance condition of the TF measurement.  The coilcraft 0805LS-273X_E 27 µH inductors series resistance was modeled as 11 Ω (as per spec sheet) and values of capacitors were those used in the actual circuit.  In the original ideal 5th order elliptical circuit there is a double dip above the corner frequency.  The series resistance of the non-ideal inductors dampens these.  I don't really want to spend much more time on mixing and matching capacitor values.  It looks like for now that the pass band ripple is acceptable and the attenuation is >40 dB at 1 MHz and 2 MHz where we are trying to block harmonic signals. I'll leave optimization of this part for now and write this off as done.

Edit Mon Feb 25 19:22:14 2019 (awade): Fixed the phase pane of the bode plot, had accidentally used magnitude and wrong units. Also fixed some spelling.

Attachment 3: Elliptical5thOrder_MeasuredVsModel.pdf
Elliptical5thOrder_MeasuredVsModel.pdf
Attachment 4: 20190222_5thOrderEllipticFilter100kHz.zip
  2297   Thu Feb 21 17:02:24 2019 awadeDailyProgressWOPOMixing down homodyne output from 1 MHz

I want to get signal from about 1 MHz down to around DC from my subtracted  homodyne photodetectors. I'm planning to do something like this:

Level 7 mixer with low pass (1.9 MHz) filter + 50 Ω terminator on output. LO will come from a
RF function generator and R port will come directly from photodetector subtraction.

This mixer has a 4.78 dB conversion loss and should do the job.  Only issue is that to operate the mixing down the low pass filter needs to be lowered from 1.9 MHz down to something pretty low to ensure the harmonic (2 MHz term is removed).  These are 4th order filters, we'd probably want the cut off to be an order of ten below the mixing frequency... 100 kHz.  I don't see this in the minicircuits catalog don't know how doable that is to make one.  I'll have a look at what the 40 m has.

Attachment 1: IMG_5106.JPG
IMG_5106.JPG
  2296   Wed Feb 20 10:38:46 2019 awadeDailyProgressWOPOFixing 532 nm polarization linearity issues

Alignment of the pumping 532 nm polarization into the WOPO is important to getting the correct phase matching condition.  For the periodically polled Lithium Niobate (LN) waveguide the phase matching is type-0: and pumping and fundamental wavelengths are in the same polarization.  The AdvR non-linear device is coupled with polarization maintaining fibers (Panda style), which are keyed at their FC/APC ends.  This means that with the correct launch polarization we should be correctly aligned with the proper crystal axis for degenerate down conversion (at the right chip temperature). 

Replacing Broadband PBS

Till now I was using non-pol maintaining patches to coupling into the WOPO fiber ends.  This should have been ok, but it is hard to figure out exactly which polarization is optimal so I switched to a pol-maintaining patch because it can be aligned separately and then the keyed connectors give you automatic alignment.  I had some issues trying to find the optimal polarization going into the fiber and I've now traced this back to the polarizing beam cubes.  I've been using Thorlabs PBS101 which is a 10x10x10 mm^3 beam cube that is supposed to be broad-band (420-680 nm).  When I checked the extinction ratio I saw Pmax=150 mW, Pmin=0.413 mW on transmission between extremes.  This is an extinction ratio of Tp:Ts = 393:1 which is much less than the spec of >1000:1.  Not sure what's going on here, the light going into the BS is coming directly from a Faraday isolator and a half-wave plate.  With some adjustment to the angle of the wave plate I can do a little better but it should be nicely linearly polarized to start with.

I've switched out the PSB101 for the laser line PBS12-1064 I remeasured extinction ratio (Pmax=150 mW, Pmin=27.6 µW) Tp:Ts = 5471:1 (better than the quoted 3000:1 spec).  This is good, at least now I know what is going on. I am also putting in an order for a 532 nm zero order quarter-wave plate, so that we can be absolutely sure we are launching in linear light always.  

Aligning light into pol-maintaining fiber

I previously thought I might be able to use the frequency modulation technique to align the light through the polarization maintaining fiber.  There is a birefringence in PM460-HP fiber of  3.5 x 10-4.  The phase between ordinary and extraordinary axes over the whole fiber length is

\Delta\phi = \frac{2 \pi \Delta n L}{c}f

Where L is fiber length, \Delta n is the birefringence and f is the laser frequency.  The idea is to launch linearly polarized light into the fiber and then at the readout place a polarizer rotated to be 90°: ramping frequency will produce an amplitude modulation on the dark fringe.  However, even with 1 GHz of frequency ramp this is only a 15 mrad effect for a 2 m fiber, its likely to be too small to see over other effects.  This is not enough to be able to fine align polarization.  

Instead I'll use the heat gun method.  I'll fire linearly polarized light into the fiber and measure the output with a crossed polarizer.  If the input polarization is correct there should be no power changes on the output as the fiber is thermally cycled. Its only two meters long so hopefully this effect is easy to see.

  2295   Tue Jan 8 11:38:49 2019 awadeLab InfrastructureGeneralRoughing pump leaking oil on bake rig

The roughing pump attached to the bake rig in the QIL (room B265B) is leaking oil.  It seems to be coming out of the box that should be filtering the exhaust (pictured below).

Its been clean up but please don't use bake rig for now and take care when entering lab.

Oil seems to be coming out of the exhast filter

I'll ask Chub to have a look.

Edit Tue Jan 8 19:58:03 2019 (awade): Chub has now fixed this. He cleaned the filter and installed a breather tube with a cloth wire-tied to the end to prevent futher spillage

Attachment 2: IMG_4862.JPG
IMG_4862.JPG
  2294   Mon Jan 7 19:27:10 2019 awadeDailyProgressHomodyneFixing homodyne fringe visibility and estimating losses

Fringe visibility of the homodyne was indeed not great.

It turns out that the combination of poor spatial overlap, polarization overlap and mode matching I couldn't see any fringes formed between the two arms of of the homodyne (when they were excited from the same laser source). 

Fixing fringe visibility

To improve spacial and polarization overlap between the two fiberized inputs I need to measure the fringe visibility, i.e. peak to peak fringe contrast. The inputs of the homodyne were connected a 50/50 fiber splitter, an extra 1 m of patch cable was used in the signal arm as follows:

Frequency modulated 1064 nm laser light is launched into both arms of homodyne
detector.  Here the signal arm patch cable is 1 m longer than the other arm giving the 
single PD FSR of 300 MHz.  The peak to peak fringes give the visablity.

With the extra length in the signal arm we have a Mach–Zehnder with an FSR  of 300 MHz.  For the Diabolo the NPRO PZT has a quoted response of 1.95 MHz/V, so 77 V is enough to see a full peak-to-peak of the fringes.

After some initial alignment of the two input beams by eye with a slow scan of the laser frequency, I was able ot see some fringes on the combined beams.  Looking at the signal on a single detector is was possible to walk the alignment between the two paths until the Vmax = 4.14 V, Vmin = 1.6 V.  Here fringe visibility is:

\textrm{Fringe Vis.} = \frac{V_\textrm{max}-V_\textrm{min}}{V_\textrm{max}+V_\textrm{min}}

With initial positioning of the fiber collimator launches I got 44 % visibility.  The quantum efficiency of the overlap scales as the square of the visibility (\eta_\textrm{h}=V^2).  Thus based on the initial placement of the fiber launches the maximum inferred quantum efficiency would be 0.19, not great.  

On closer inspection of the interference pattern, it was apparent that the interference was forming a bull's eye pattern.  The propagation​ distance from the signal arm was 20 mm shorter than that of the LO arm.  I moved the fiber launch back to match the arm lengths and realigned the beams.  Now with 125 mm distance for both arms to the beam splitter I was able to optimize to Vmax = 4.9 V and Vmin = 0.320 V, giving a visibility of 88%. That means a minimum quantum efficiency due to overlap of 0.77.  This could be an underestimate of the visibility and efficiency if the polarizations were not optimal.  More likely the power ballance between the arms would have biased the measurement.  I didn't note down the power measured from each of the launches, from memory it was to within 5%.

Another thing to note is that the patch cables used in the launch are slightly different.  Even though the collimators are both F240APC-1064 the slightly different MFD of the two fibers means that the launched beam will have a slightly different waist (position​ and radius).  Maybe mode matching could be improved with the purchase of a matching patch for the signal path to the LO path.

Total estimates of losses (Worst case)

Total estimated losses are max 0.2 dB/cm within the waveguide (total max .91), 0.5 at the fiber WOPO butt coupled interface, 0.7 dB (typical/max?) insertion loss at fiber to fiber interface x1 (~ 0.15), 0.23 due to mode overlap loss at HD, 0.15 at the photodiode themselves.  This would total 0.25 quantum efficiency​.  From this the lower bound of max squeezing would be

 SQZ_\textrm{Max} = 10 \log[1-\eta_\textrm{total}]

This would put the best case squeezing at about -1.25 dB, given all the above estimated losses.  

Attachment 1: scan_19-01-08_20-07-14.pdf
scan_19-01-08_20-07-14.pdf
  2293   Fri Jan 4 19:49:51 2019 awadeDailyProgressHomodyneTweek up fiber alignment and reboot of homodyne

[From before weekend]

Its not clear that in the past I had good mode overlap in the homodyne between the signal port and the local oscillator (LO).  Today I looked into that and found that some of the signal light is clipping the west PD, which would indicate that the alignment is poor.  

I used a fiber splitter to launch 1064 nm out of both the LO and signal ports from the same laser source.  Looking at the beam over lap in the near field and far field, it is pretty obvious that the beams are not coincident. They clearly hit different points on the card, not great.  

The way I configured this was to connect one leg of the 50/50 fiber splitter to the launch of the signal path and the other into the existing  2m + fiber paddle in the LO path. We should expect to see some fringing as the laser drifts or small temperature changes cause some LF drift (in this case FSR = 150 MHz). With a little tweeking of alignment, without moving any opitics I think I have them overlapping but don't see coherent fringing.  Tomorrow I'll need to remove a steering mirror.  Also problem could be caused by bad polarization, that is another thing to check.  

One way check goodness of alignment would be to drive the laser frequency with waveform much faster than thermal drift and walk the alignment of the LO + signal until the fringing peak to peak is maximized.  For the Diabolo, the PZT response is 1.95 MHz/V, 77 V pp should be enough for a full fringe of this 2 m mismatch MZ.  The only question is wether the fiber launch from the WOPO end is going to be seated exactly the same between fiber insersions and whether the slighly different MFD of the fiber will mess with the MM.

More to come.

Quote:

Today I restarted the Diabolo.  Still good on the high power generation, so that's good.

Checked the throughput of the 532 nm fiber.  Looks like from 2.04 mW at the input coliumator I get 1.05 mW at the output. This makes it 51%, ok, enough to work with.

Also checked 1064 nm alignment and MM into fiber.  There from 6.0 mW I get about 2.8 mW out from the other side.  Not the best but more than enough given that I think I only need about 2mW at most.

---

I also resurrected the homodyne. The alignment on PDs was checked.  The polarization was an arbitrary elliptical state (beam splitter is supposed to be optimal for s-pol).  I think this was left in this state because I was doing some other optimization before where I was balancing​ power using polarization.  It could also be the case that the paddle polarization controller has undergone temperature drift.  I used a beam cube to make sure that there is now only linearly polarized in the vertical direction.

With the electronic HD ballance trimmer I was able to bring the homodyne common mode rejection ratio (CMMR) down to 65 dB. This means that there is a 65 dB suppression of noise induced from RIN in the optical local oscillator.  I fine tuned this by injecting a 50 mVpp @ 3.1 kHz sine into the laser diode current modulation into (0.1A/V) on the back of the Diabolo controller and comparing the transfer function with one PD blocked and then both un-blocked.

It looks good to have a look at the impact of PZT scan on subtracted signal tomorrow.  I also need to check how much 1064 nm seed light I I can expect to get through to the WOPO copropagating with the 532 nm pump. 

 

Attachment 1: 20190104_TopViewSQZHDSetup.jpg
20190104_TopViewSQZHDSetup.jpg
  2292   Wed Jan 2 18:51:53 2019 awadeDailyProgressHomodyneTweek up fiber alignment and reboot of homodyne

Today I restarted the Diabolo.  Still good on the high power generation, so that's good.

Checked the throughput of the 532 nm fiber.  Looks like from 2.04 mW at the input coliumator I get 1.05 mW at the output. This makes it 51%, ok, enough to work with.

Also checked 1064 nm alignment and MM into fiber.  There from 6.0 mW I get about 2.8 mW out from the other side.  Not the best but more than enough given that I think I only need about 2mW at most.

---

I also resurrected the homodyne. The alignment on PDs was checked.  The polarization was an arbitrary elliptical state (beam splitter is supposed to be optimal for s-pol).  I think this was left in this state because I was doing some other optimization before where I was balancing​ power using polarization.  It could also be the case that the paddle polarization controller has undergone temperature drift.  I used a beam cube to make sure that there is now only linearly polarized in the vertical direction.

With the electronic HD ballance trimmer I was able to bring the homodyne common mode rejection ratio (CMMR) down to 65 dB. This means that there is a 65 dB suppression of noise induced from RIN in the optical local oscillator.  I fine tuned this by injecting a 50 mVpp @ 3.1 kHz sine into the laser diode current modulation into (0.1A/V) on the back of the Diabolo controller and comparing the transfer function with one PD blocked and then both un-blocked.

It looks good to have a look at the impact of PZT scan on subtracted signal tomorrow.  I also need to check how much 1064 nm seed light I I can expect to get through to the WOPO copropagating with the 532 nm pump. 

  2291   Mon Dec 24 16:46:41 2018 awadeDailyProgressWOPOMM into 532 nm patch fiber (to be hooked up to WOPO)

I had another look at MM the 532 nm light into the P3-460B-FC-2 patch cable.  After walking some lens positions and mirror pointing based on ATF:2257 I found that I could get 1.07 mW output from launch of 2.3 mW (46.5 %) through the fiber.  I expect the most I'll need is about 30 mW at the output which will put the amount of required power at about 64 mW.  There is now enough 532 nm laser light to do this and I think its within the tolerances of the fiber to have that much launched in.

One thing I did notice, on closer inspection of the fiber ends, is that there is a little bit of damage on one end (labeled A) of the P3-460B-FC-2 patch.  This is pictured below.

Yellow patch SM fiber (P3-460B-FC-2) damage at end A.

I tried a bit of gentle cleaning with some fiber cleaning cloth (Thorlabs FCC-7020)  but this appears to be a burn mark when zoomed in. Sorry doesn't capture very clear on my phone through the fiber microscope. I swapped the end that was in the fiber collimator (F240APC-532) but didn't increase the fiber launch efficiency​.  Not sure when this got damaged, but I might have exposed it to excessive power at some stage. 

I still think that the fiber through put should be sufficient, if it turns out to be a problem then re-ording shouldn't be an issue.  Thorlabs usually seems to make these fiber items next day.

Next steps

Next step is to tune up the 1064 nm alignment into the LO launch fiber and see if I can get the new PZT mounted mirror to scan as expected.  Need to figure out if there is a way to calibrate the phase on this so that I can check it is actually scanning phase (the last PZT broke and I had no idea).

Need to figure out how much 1064 nm light can be co-propagated in the patch waveguide.  The previous dismal power values might be because I made very little effort to mode match,  I'll look at what lenses I have available and see if I can boost the power through put to something respectable.

  2290   Mon Dec 24 12:07:18 2018 KojiDailyProgressWOPODiabolo: back to power with deliberate pre-misalignment

Great recovery job!

 

  2289   Mon Dec 24 10:08:49 2018 awadeDailyProgressWOPODiabolo: back to power with deliberate pre-misalignment

I think I've managed to bring the power of the Diabolo back to a usable level for WOPO.

More power with misalignment

I had another look at the alignment into the SHG.  This time I systematically trialed a range of cavity misalignments to deliberately set the pointing into the cavity to be wrong for the fast scanning mode of the control unit (cold) but to allow for the locked (hot) cavity to expand into correct alignment.  This video shows the SHG output as the cavity is locked IMG_4766.MOV,  this seems to show the expansion affecting mostly the horizontal axis. However, eventually I found that the locked cavity had a preferred misalignment on the vertical axis (in terms of the nobs that I needed to turn).  In reality its a mixture of horizontal and vertical once all the DOF are traced through the lenses and mirrors.

I was able to get a steady 160 mW of power with this new misalignment/alignment strategy.  The photos below shows the 1064 nm transmission peaks and error signal for a combination of (mostly) vertical and (a little) horizontal misalignment​ that gave greater power. 

SHG transmission peaks for alignment that delivered 160 mW
PDH error signal corresponding to above misalignment (202 mVpp amplitude)
Steady state power, looks like it peaks at much higher power
(823 mW, but couldn't find a optimal alignment and
oven temperature point that exceeded about 165 mW)

The oven operating temperature was at 110.14 C with a laser diode current of 2.102 A.  PDH gain was set to 0.56 and offset was set to 5.1.

Upping the oven temperature limit

The Diabolo SHG oven set point temperature range doesn't put the ideal phase matching temperature at the center of the available values.  In fact, the optimal power for the above-mentioned power boost was achieved pretty much of the edge of the available range (top of range is 110.26).  I popped the lid on the controller box and traced through all the PCBs to find out how the temperature range was trimmed. The front panel potentiometer (1k) is trimmed with a small set screw (blue) potentiometer located directly to the low left of the temperature knob (when viewed from the component mounting side).  I trimmed the set point temperature value so that the upper value was 111.00 C, giving an extra 0.75 C of headroom above the ideal phase matching set point.  

I found that at the very upper edge of the new range the thermal control loop started to loose some stability.  I limited the increase of temperature to the 111.00 C point and will use some caution when adjusting temperatures at these upper ranges to make sure I don't get oscillations in power.

 

 

 

 

Quote:

The strategy locking at lower laser power and optimizing phase matching temperature by slowly walking the temperature and laser power up won't work: the autolocker mode of the Diabolo SHG won't engage at lower laser power. Instead I tried a brute force approach of stepping temperature and re-locking the cavity in 0.05 C increments of set point temperature.  I've plotted the result below.  Error bars are based on standard deviation on 32 averaged measurements over 10 seconds.

I stopped at 106.00 C as  power output was not improving beyond that point. Maximum temperature is clipped at the controller at 110.26 C. Peak output of 66 mW is actually  at 110.20 C. This is counterintuitively higher than the previous set point but might be good enough to work with, as long as its stable over longer periods.

Still not clear what has changed to shift the ideal SHG over set point to 1.4 C higher than it was before.  Maybe the SHG oven needs to go even a little hotter, but we can only access tempertures up to 110.26 C. I'll take what I can get.

Some other laser settings: 

  • Injection current: 2.077 A
  • Laser crystal temperature: 26.32 C
  • SHG PDH gain: 0.5
  • PDH offset: 4.95
  • Noise eater: off

Data and plotting notebook are attached in a zip below: 20181220_SHGLockedPowerVsSetPointTemp.zip

 

  2288   Fri Dec 21 16:14:42 2018 RahulDailyProgressCryo vacuum chamberPump down results

I switched on the pumping station this morning at 10 am and roughly after 6 hours of pumping the pressure in the vacuum chamber is 8.2*10^-6 Torr. The pressure is being monitored using two gauges (wide range gauge and Ion gauge). However due to some reason the gauge controller is unable to switch on the Ion gauge (there is an inbuilt safety in the controller to protect the low pressure gauge, so this could be one of the reason).

 

I will keep the pumps running overnight and through out the weekend to monitor the pressure.

Attachment 1: Pressure.jpg
Pressure.jpg
Attachment 2: PLot1_pressure.png
PLot1_pressure.png
  2287   Fri Dec 21 15:51:27 2018 awadeDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

No I don't think so.  The temperature of the oven is important for the phase matching condition.  Its the nob you can turn so that the refractive index for both 532 nm and 1064 nm is just right so that both traveling waves stay in phase as they propagate in the non-linear crystal.  Otherwise the phase precesses in one wavelength relative to the other causing repeated amplification and de-amplification as the light propagates (rather than just amplification when they are in phase).  

Its true that temperature change induces expansion​ and dn/dt changes that will change the round trip effective length/phase. However, the PZT in the SHG should pick up the slack of from changes in the crystal round trip phase due to temperature.  We can tune the laser temperature to bring the PZT to the center of its range but at this stage it doesn't look like its railing the range of this actuator.  

My guess is that self heating affects the relative alignment of the crystal HR surface relative to the PZT mounted (curved) mirror, this small change in alignment will shift the eigen axis of the cavity and mean that the locked hot cavity will have a slightly different optimal alignment to the unlocked case.  Its much harder to walk alignment of 1064 nm while keeping cavity locked.

 

Quote:

This means that you want to make the SHG crystal longer. Is that true? If so, can you change the temperature for the optimal phase matching by tuning the 1064 crystal temperrature? I suspect you need to cool the YAG crystal, but I am not sure what is the thero-optic constant of the SHG crystal, and how much you can gain from this.

 

  2286   Fri Dec 21 15:33:26 2018 KojiDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

This means that you want to make the SHG crystal longer. Is that true? If so, can you change the temperature for the optimal phase matching by tuning the 1064 crystal temperrature? I suspect you need to cool the YAG crystal, but I am not sure what is the thero-optic constant of the SHG crystal, and how much you can gain from this.

  2285   Fri Dec 21 15:23:02 2018 RahulDailyProgressCryo vacuum chamberChamber closed for 1st pump down test

The cryo vacuum chamber was closed yesterday (Thursday) for the 1st pump down test. The figure below shows various components which was attached on to the CF flanges. Later I will post the results of the ongoing pump down test.

Attachment 1: Picture1.jpg
Picture1.jpg
Attachment 2: Picture2.jpg
Picture2.jpg
Attachment 3: Picture3.png
Picture3.png
Attachment 4: Picture4.jpg
Picture4.jpg
  2284   Thu Dec 20 19:20:50 2018 awadeDailyProgressWOPODiabolo: trialing SHG oven set temperature points to see if there is a maxima when locked

The strategy locking at lower laser power and optimizing phase matching temperature by slowly walking the temperature and laser power up won't work: the autolocker mode of the Diabolo SHG won't engage at lower laser power. Instead I tried a brute force approach of stepping temperature and re-locking the cavity in 0.05 C increments of set point temperature.  I've plotted the result below.  Error bars are based on standard deviation on 32 averaged measurements over 10 seconds.

I stopped at 106.00 C as  power output was not improving beyond that point. Maximum temperature is clipped at the controller at 110.26 C. Peak output of 66 mW is actually  at 110.20 C. This is counterintuitively higher than the previous set point but might be good enough to work with, as long as its stable over longer periods.

Still not clear what has changed to shift the ideal SHG over set point to 1.4 C higher than it was before.  Maybe the SHG oven needs to go even a little hotter, but we can only access tempertures up to 110.26 C. I'll take what I can get.

Some other laser settings: 

  • Injection current: 2.077 A
  • Laser crystal temperature: 26.32 C
  • SHG PDH gain: 0.5
  • PDH offset: 4.95
  • Noise eater: off

Data and plotting notebook are attached in a zip below: 20181220_SHGLockedPowerVsSetPointTemp.zip

Attachment 1: plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf
plot20181220_SHGLockedOutputVsOvenSetpointTemperature.pdf
Attachment 2: 20181220_SHGLockedPowerVsSetPointTemp.zip
  2283   Wed Dec 19 16:48:18 2018 awadeDailyProgressWOPODiabolo: baseline numbers for coherent

Contact with Coherent re Diabolo

I got in contact with Coherent they have provided some instructions on aligning the laser cavity, something I've already attempted and knew how to do.  For future reference I've attached the instructions on the wiki (they are confidential): they can be found attached HERE.

More details on Diabolo manual and documentation can be found on the Manuals-And-Datasheets ATF wiki page.  

Temperature tuning

I've been attempting to find sweet spot for phase matching temperature between oven and self heating.  The FWHM of the phase matching temperature of the SHG lithium niobate crystal appears to be on the order of about 0.2 C which makes it quite narrow.  Given there is quite a bit of self heating in the cavity when it is locked up, it would be very easy to miss the correct temperature when setting the crystal's oven set point.

Koji/rana has suggested turning down the laser power, locking and then gradually tuning the oven temperature to get to ideal set point as we increase power.  I've attempted this and found that the autolock requires some minimum amount of laser power to engage.  I opened up the controller box and traced a few things through the PCBs, but it is not immediately apparent which pot adjusts the autotune threshold value for power.  I've emailed the contact at Coherent and they are making inquires to see if they can tell me what to adjust to lower this threshold manually.  For now I will attempt a few more full power locks at carefully spaced values of crystal temperature to see if I am able to happen upon a sweet spot.

  2282   Mon Dec 17 13:52:58 2018 AidanLab InfrastructureCleanlinessOil spill on North table

I discovered that there is an oil (or solvent) spill on the north (cryo chamber) table. This is over the IQIM breadboard and over the optical table itself. I shall engage cleaning procedures.

 

Attachment 1: IMG_7827.jpg
IMG_7827.jpg
Attachment 2: IMG_7828.jpg
IMG_7828.jpg
  2281   Thu Dec 13 16:24:06 2018 awadeDailyProgressWOPODiabolo: baseline numbers for coherent

Some numbers for the Diabolo laser for reference (to quote to Coherent, if they get back to us with help).

Diabolo numbers (Dec 13, 2018)
Parameter Value
Fundamental (1064 nm) laser power exiting laser from tap off BS 372 ± 2 mW
Fundamental (1064 nm) incident on SHG (measured after steering mirror 1) 1.32 ± 0.02 W
Total fundamental (summed from above two mesurements) 1.692 ± 0.02W
SHG cavity scan TEM00 peak (voltage trans diode) 1.21 V (1 MΩ imp oscilloscope)
SHG cavity scan second HOM peak 38 mV (1 MΩ imp)
SHG cavity scan third HOM peak 17 mV (1 MΩ imp)

 

To characterize the phase matching temperature I fast scanned the SHG and measured the peak 532 nm output for a range of temperatures in 0.05 C steps.  Acoss a single sweep of the cavity resonances there is some change in the peak 532 nm output.  This change accross the sweep range is likely an alignment thing with the shape of the PZT sweep path.  Plot below shows the peak 532 nm output power measured on a Thorlabs PD (PDA100A, 0 dB gain with 1.3 ND filter in front). Vertical axis units are less relivant than the shape and peak possition of the curve.  Error bars were estimated based on spread of peak amplitudes (the largest error in the mearument).

We see a peak conversion at 109.8 C with a FWHM of about 0.2 C

I've attached data and python notebook for plotting in a zip below.

 

Quote:

I've contacted the people at Coherent about getting instructions for repair or whether they can still service this unit as I don't want to waste a lot of futile time on trying to optimize a unit that might be broken/degraded.

The technicians I spoke to didn't seem to know a lot about the Diabolo unit but will forward on to their manufacture for further information/assistance.

 

Attachment 1: plot20181218_PhaseMatchingCurveFastScanDiaboloSHG.pdf
plot20181218_PhaseMatchingCurveFastScanDiaboloSHG.pdf
Attachment 2: 20181213_SHGScanWithLNTempTuning.zip
  2280   Thu Dec 13 14:01:13 2018 awadeDailyProgressWOPODiabolo: contacted coherent

I've contacted the people at Coherent about getting instructions for repair or whether they can still service this unit as I don't want to waste a lot of futile time on trying to optimize a unit that might be broken/degraded.

The technicians I spoke to didn't seem to know a lot about the Diabolo unit but will forward on to their manufacture for further information/assistance.

 

Quote:

I've managed to vastly improve the power output of the SHG by doing a little cleaning inside the box.

After some consideration I decided that the only remaining possibility was something wrong inside the SHG. Its a hermetically sealed box so I cleaned around the area (to keep dust down) and then removed the lid.  Some of the screws were loose, so its likely that somebody has opened this before.  Below are a few pictures of the inside of the Diabolo SHG. 

 

 

  

Looking through an IR viewer I could see that there was a large absorber on the front face of the crystal.  I didn't get a picture but it was dead center and scattering a lot of light given the tightness of the beam. There was also other smaller scatters around the edge of the front surface. An clean N2 ionizing air gun was used to blow in and around the cavity and its electronics. There was a lot of dust blown out from inside the unit; the SHG has evidently been opened at some point in a dirty lab.  The spot in the middle of the front face of the SHG came loose with very gentle bushing with a corner of a lens cleaning tissue (MC-5).  I thoroughly flushed the inside of the unit again with 40 psi clean N2 and then replaced the lid.  Just before sealing the box I cracked the lid and flushed again with the N2 gun to fill the box with clean dry air.  The screws were sealed down with 2.25 Nm of torque.

I ramped the laser back on and after some small small alignment tweaks I was able to get an initial large boost in 532 nm laser output but it quickly dropped to the lower ~12 mW output again. It looks like, after lock, the output light goes through a few fringes and then either settles on a HOM or a weaker fundamental mode.  My guess is that in the locked state the crystal/mirror is heating up and mudslinging the cavity.  I don't really want to pull the unit apart into is constituent pieces to inspect.  I might need to have a closer look at the crystal front face absorber and the state of the input coupler mirror.  Over time some non-linear crystals do develop either grey tracking or permanent damage along the tighly focused points of 532 nm light.  It could be that the unit has gone past a tipping point but this is hard to see from just peaking in from the top of the unit.

Not sure what to do next.

---

Side note: Its seems very hard to make any alignment changes from the inside: my guess is that alignment is done from the three outside exposed hex screws on the front coupler. Its unlikely with such a short cavity much alignment changes would be needed.  For now its best to leave it as it is.

 

  2279   Wed Dec 12 14:40:17 2018 RahulDailyProgressCryo vacuum chambercleaning of shields and cold plate

The radiation shields and the cold plate have been cleaned (class B) and baked and can be used in vacuum. I brought them back to the lab and will ship the outer shields for Gold plating.

Attachment 1: IMG_0238.jpg
IMG_0238.jpg
Attachment 2: IMG_0237.jpg
IMG_0237.jpg
  2278   Tue Dec 11 16:26:07 2018 RahulDailyProgressCryo vacuum chambercleaning of shields and cold plate

The radiation shields and the cold plate have been cleaned twice at the bath at 40m. They are currently undergoing baking for 12 hours, will pick them up on Wed morning once finished.

  2277   Fri Dec 7 20:31:06 2018 awadeDailyProgressWOPODiabolo: cleaning dust out of SHG unit, self heating issues

I've managed to vastly improve the power output of the SHG by doing a little cleaning inside the box.

After some consideration I decided that the only remaining possibility was something wrong inside the SHG. Its a hermetically sealed box so I cleaned around the area (to keep dust down) and then removed the lid.  Some of the screws were loose, so its likely that somebody has opened this before.  Below are a few pictures of the inside of the Diabolo SHG. 

 

 

  

Looking through an IR viewer I could see that there was a large absorber on the front face of the crystal.  I didn't get a picture but it was dead center and scattering a lot of light given the tightness of the beam. There was also other smaller scatters around the edge of the front surface. An clean N2 ionizing air gun was used to blow in and around the cavity and its electronics. There was a lot of dust blown out from inside the unit; the SHG has evidently been opened at some point in a dirty lab.  The spot in the middle of the front face of the SHG came loose with very gentle bushing with a corner of a lens cleaning tissue (MC-5).  I thoroughly flushed the inside of the unit again with 40 psi clean N2 and then replaced the lid.  Just before sealing the box I cracked the lid and flushed again with the N2 gun to fill the box with clean dry air.  The screws were sealed down with 2.25 Nm of torque.

I ramped the laser back on and after some small small alignment tweaks I was able to get an initial large boost in 532 nm laser output but it quickly dropped to the lower ~12 mW output again. It looks like, after lock, the output light goes through a few fringes and then either settles on a HOM or a weaker fundamental mode.  My guess is that in the locked state the crystal/mirror is heating up and mudslinging the cavity.  I don't really want to pull the unit apart into is constituent pieces to inspect.  I might need to have a closer look at the crystal front face absorber and the state of the input coupler mirror.  Over time some non-linear crystals do develop either grey tracking or permanent damage along the tighly focused points of 532 nm light.  It could be that the unit has gone past a tipping point but this is hard to see from just peaking in from the top of the unit.

Not sure what to do next.

---

Side note: Its seems very hard to make any alignment changes from the inside: my guess is that alignment is done from the three outside exposed hex screws on the front coupler. Its unlikely with such a short cavity much alignment changes would be needed.  For now its best to leave it as it is.

Attachment 5: IMG_4766.MOV
  2276   Fri Dec 7 16:08:45 2018 RahulDailyProgressCryo vacuum chamberRadiation shields

The cold plate and radiations shields (100K, 50K and bottom 100K shields) are at 40m for cleaning and baking. The outer shields will be cleaned and then sent for gold plating. The inner shield and cold plate be baked after cleaning. 

Attachment 1: shields_1.jpg
shields_1.jpg
Attachment 2: shield_2.jpg
shield_2.jpg
Attachment 3: shields_bottom.jpg
shields_bottom.jpg
  2275   Fri Dec 7 14:43:20 2018 RahulLab InfrastructureGeneralCardboard and packaging in the lab

I have removed all the cardboard boxes and packaging materials which belongs to cryo-vacuum chamber.

edit awade Fri Dec 7 20:30:03 2018: Awesome thanks so much.

Quote:

There are a bunch of cardboard boxes and paper packing material in the QIL lab and the dust levels are at the point that I can see particles despositing on all surfaces.

Can we please make an effort to remove all these paper materials pronto from the lab and get a weekly wet floor sweep going? 

Cardboard materials should removed from shipped items before they make it into the lab.

We don't have a air particle counter in this lab, maybe we should.

 

  2274   Fri Dec 7 11:59:29 2018 awadeLab InfrastructureGeneralCardboard and packaging in the lab

There are a bunch of cardboard boxes and paper packing material in the QIL lab and the dust levels are at the point that I can see particles despositing on all surfaces.

Can we please make an effort to remove all these paper materials pronto from the lab and get a weekly wet floor sweep going? 

Cardboard materials should removed from shipped items before they make it into the lab.

We don't have a air particle counter in this lab, maybe we should.

  2273   Thu Dec 6 20:04:50 2018 awadeDailyProgressWOPODiabolo 532 nm power drop and retuning

I've continued to have problems getting the Diabolo laser SHG to lock. 

I get nice clean fundamental resonance peaks when fast scanning the SHG cavity length and optimizing the 1064 nm alignment into the SHG . However, when I switch to engage I see an initial power jump before the power dips and then settles to what looks like a higher order mode.  It seems like the heating of the NL crystal when locked is misaligning the eigen-mode of the cavity so that the cold scan of the cavity is optimal until locking and the higher circulating power heats things up.  

I tried a different approach, slow scanning close about the fundamental (to keep things warm) and attempted to maximize the fundamental peak.   This was done by narrowing the scan range and then tuning the laser frequency (using NPRO crystal temp) to be close to resonance.  This improved the power output by a few tens of percent to ~27 mW, but is still well short of the previous performance.  This current power output is unworkable because we need at least 60 mW to overcome losses coupling into the WOPO and also because it is not stable in power and drifts around a bit and sometimes unlocks (probably because it is on the edge of some thermal state of the system).

I also tried tuning 1064 nm alignment into the cavity with the SHG locked.  This is tricky because that for each adjustment it takes a little time for the thermal state of the crystal to stabilize and settle. It improved power output maybe 10% but wasn't clearly leading to good operating point quickly.

Fixing this

So it seems the unit was previously in a combination of states that worked once the SHG was locked.  Its been hard to reproduce that operating point. Its possible that since 2004, when the unit was manufactured, that there has been some degradation to the crystal.  Also, maybe the reason it has failed to perform to spec might be that the SHG cavity itself is slightly mialigned.  I'm reluctant to open up the SHG unit as it is hermetically sealed.  I assume they flood it with N2 to keep everything in a dry no oxygen controled environment but at this stage maybe it doesn't matter.  Tweaking the input coupler on the SHG will be super sensitive so would be a delecate task, but I'm at a loss as to how to find an optimal operating point to return to normal operation.

 

  2272   Thu Dec 6 16:49:40 2018 RahulDailyProgressCryo vacuum chamberChamber and bottom plate assembly

The bottom plate and the vacuum chamber (collar) has been assembled – thanks to Aaron for helping me out. The O-ring holds in well in the dovetail groove and makes it really easy (without falling off) to assemble the two components together.

Attachment 1: O-ring_bottom.jpg
O-ring_bottom.jpg
Attachment 2: Bottom_collar_assembly.jpg
Bottom_collar_assembly.jpg
  2271   Thu Dec 6 12:44:18 2018 RahulDailyProgressCryo vacuum chamberFabricated parts arrived

Cold plate and the radiation shields have arrived. I will perform a mock assembly of these fabricated parts to check if they are as per our drawings. Next, the outer radiation shields will be shipped for gold plating the outer surface. Rest of the fabricated parts will be cleaned (lots of grease and finger marks visible) and then baked.

Attachment 1: Fabricated_Cold_plate.jpg
Fabricated_Cold_plate.jpg
  2270   Wed Dec 5 16:49:50 2018 RahulDailyProgressCryo vacuum chamberSchematic for the flanges

The vacuum chamber has 12 flanges (four 4-5/8 and eight 2-3/4 flange size). For the pump down test I will attach the pumping station to the 4-5/8 flange and close down the other three flanges with blank flange. Similarly out of the eight smaller flanges (2-3/4 size) five will have Tees (for 2 gauges), up to air valve, two optical windows, cryo-feedthrough and the remaining with blank flanges (or else I can change them with other feedthroughs to test them out for vacuum leakage). Just for information - the pumping station is attached to a reducer which will house the rotary valve. Reducer is requred to match the flange size of the pumps (4-51/2 siz - CF63 size) and the vacuum chamber (4-5/8 size). The rotary valve will be attached to the hose which will then be attached to the flange of the vacuum chamber.

Attachment 1: Schematic_chamber.jpg
Schematic_chamber.jpg
  2269   Wed Dec 5 16:18:50 2018 RahulDailyProgressCryo vacuum chamberbottom plate

Aaron and Rahul - The bottom plate of the chamber was upside down (Gabriele and I did it as we were trying out few things during assembly last week). Now, the bottom plate has been inverted (using crane and trolley) and is ready for the chamber to be brought in. I am giving the chamber and the plates a thorough wipe (speckles of dust easily visible) and the grooves for the O-ring also looks slightly dusty. Probably I should get clean plastic sheets to cover the vessel when the top lid is open.

  2268   Tue Dec 4 19:34:22 2018 awadeDailyProgressWOPODiabolo 532 nm power drop and retuning

I restarted the Diabolo laser after it had been shut down for a lab tour.  The using the throlabs S130C head the 1064 nm was 355 mW which is about what we expect (previously it was measured to be 302.3 mW @ 1064 nm. see ATF:2103).  However I was unable to recover the 532 nm full power from the SHG; power from the SHG was down to 5.9 mW (previously ~300 mW).

It seems like mode of the 532 nm is not that clean.  Its likely that the PDH servo is not locking to the optimal mode any more. Because the locked cavity self heats with the high circulating power, often the steady state operation the SHG has a different lock point from the cold start.  Some days I kind of scan it for a bit, lock on the best mode and then come back after a few minutes and relock on the brightest mode.

The last record I made of the Diabolo and its SHG settings was in ATF:2103.  It seems like the doubling crystal temperature has been changed at some stage from 102.87 C to 108.84 C.  We could have done this with Dhruva but it might not have been noted down.  In this original post I noted down values and the SHG was set to its lower bound, this didn't match up with what was in the manual that I had at the time.  Later Koji found the correct manual for this unit (see ATF:2104).  It seems that the manufactures value for peak phase matching was 109.20 C.

Snapshot of settings

Before I mess up the settings by tuning some nobs, here is the present settings for future reference:

Laser diode A temp = 19.69 C
Laser diode B temp = 20.32 C
Injection current  = 2.1 A
Laser crystal temp  = 23.42 C

SHG unit settings were
Double crystal temp = 108.83 C
Offset = 5.09
Gain  = 0.4
Scan amplitude  = 0 (you turn this up to scan SHG cavity)

Fixing alignment of 1064 nm into SHG

I then scanned the SHG cavity to see the transmission peaks and PDH error signal (BNC outputs are on the back of the control unit).  It looked about the same as that reported in ATF:2103.  I then popped the lid on the laser head and tuned the alignment of the two steering mirrors into the SHG to maximize the dominant mode while minimizing the secondary peaks. 

Inside of Diabolo laser head unit.  Steering mirrors were adjusted to optimize alignment.

Trace is displayed below (second attachment).  The lopsidedness of the main peaks are heating effects in the non-linear crystal.  Compare to ATF:2103, ratio of secondary peaks to primary has gone from ~0.25 to <0.05. This should give me a much better power output when relocked.

Still no full power on lock

After replacing lid on laser and attempting to relock the SHG unit I still get the weak mode.  It doesn't look like a HOM but its hard to tell.  

Possible issues are: that we might be getting bad phase matching temperature after locking; self heating is pulling pulling SHG cavity mode to saturate the PZT range; and maybe during the power cycle of the SHG unit, knobs were changed that have messed up the stability of the PDH loop.

I checked the error signal.  Its lopsided, like the transmission peaks, because of the self heating of the Lithium Niobate crystal as we cross resonance. However, zooming in on the oscilloscope seems to show the peak corresponding to the zero crossing point of the error signal. I also played around with the loop gain to find that it is nicely below the point at which it starts ringing.  Its a bit hard to get get an openloop transfer function without getting into the guts of the control box.

SHG PDH signal x10 along with transmission peaks.
Lopsidedness is due to Lithium Niobate self heating, otherwise it looks healthy, if not a bit small.

Bad locking of the SHG is still unresolved.  I will do more troubleshooting today as I can't proceed without 532 nm light.

Attachment 2: plot20181204_DiaboloSHGRaligned1064nm.pdf
plot20181204_DiaboloSHGRaligned1064nm.pdf
Attachment 3: plot20181204_DiaboloSHGRErrSig.pdf
plot20181204_DiaboloSHGRErrSig.pdf
Attachment 4: 20181204_RetuneDiaboloSHGAlignment.zip
  2267   Tue Dec 4 15:56:29 2018 RahulDailyProgressCryo vacuum chamberO-ring which fits

The viton O-ring shipped by Nor-Cal fits the chamber groove, given below is the correct specification for future reference. I am buying few more for spares.

2-473 Viton , ID 23.940'' OD 24.490'' Width 1/4''

Attachment 1: IMG_0170.jpg
IMG_0170.jpg
Attachment 2: IMG_0171.jpg
IMG_0171.jpg
  2266   Thu Nov 29 17:00:37 2018 RahulMisc Nor-Cal: leak rate of the chamber (collar)

Nor-Cal indicated leak rate of the collar is 9.8*10^-10 std. cc/ss (Std.cc/sec= One cubic centimeter of gas flow per second at 14.7 psi of pressure and a temperature of 77 F.). I have uplaoded their spec. sheet (which I hope is their measured data) for future reference. The surface finish is Electropolish.

THIS IS THE ELOG POLICE
DON'T PUT DATA SHEETS OR ANYTHING PROPRIETARY ON OUR PUBLIC
 WEB SITE
Put it on a wiki page and put the link here. Thanks! (Koji)
 

I apologise for my mistake, given below is the wiki link for the outgassing rate data supplied by Nor-Cal for our chamber.

https://nodus.ligo.caltech.edu:30889/ATFWiki/lib/exe/fetch.php?cache=&media=main:experiments:cryovacchamber:outgassing_norcal.jpg

  2265   Thu Nov 29 16:52:50 2018 RahulSummary Cryo-vacuum chamber

I would like to give a brief update on the ongoing effort for the assembly of the cry-vacuum chamber at the QIL. Gabriele has been kind enough to spare some time to help me with safe crane operation, considering the chamber and plates are 70 kg each.

At first, I prepared an aluminum spacer (made out of bosch extrusion) which is 90 mm in height. The idea is to lift the bottom plate (which has 16 through holes and counterbore for bolting purpose) such that the bolts can be easily inserted from the bottom. The spacer is strong enough to take the load of the entire assembly (i.e. around 200 kg, bottom and top plate + collar). The assembly will be resting on this spacer, even during the in-situ baking procedure. After baking and pump down testing is complete, the spacer will be removed and the vacuum chamber will be resting flat on the optical bench.

Next, with the help of a crane we tried lifting the collar using eye bolts and nylon slings, however the suspension point was way too high giving us no room for the crane to lift. This required shortening the length of the nylon sling. Firstly, I got shorter length eye bolts (2.5 inch against 4 inch) and secondly I used carabiners to tie up the sling. Using this set up we were able to lift the collar successfully.

The collar has groove for single dovetail O-ring (pictured attached) at the top and bottom surface.

The specifications for the viton O-ring was provided by Nor-Cal. I bought the viton O-ring (2-474V, 24.94’’ ID) from Kurt Lesker, the dimensions of which is given below,

O-RING,FKM,24.940"ID X .275" W(ACT),25"ID X 1/4"W (NOM),ISO 630

The challenge is to keep the O-ring held in the groove at the bottom of the collar (against gravity) during it’s mating with bottom plate. However, the O-ring always falls off. At this point I am not inserting the O-ring completely inside the groove (and I will come to this point next) and just slightly pushing it in. I want the O-ring to move in naturally from the pressure created by the plate and collar. I contacted Zach since he has faced similar issues with his tank (Chris gave me this information) and he used vacuum grease to keep the O-ring sticking on to the groove. However, outgassing from the grease could be a point of concern.

Next, I tried to insert the O-ring completely inside the grove, strangely in this case I find it to be longer by at least an inch. The inner diameter of the groove is 25 inches.

I contacted Nor-Cal to cross-check the size they have recommended. They realized that they have made a mistake with the size and have agreed to ship one size smaller O-ring (i.e. 2-473 having an ID of 23.94) which should give a stretch of 4%. I will use this once we receive it.

I found a Youtube video of the O-ring installation in a half or full dovetail groove. Given below is the link, since this is different from the usual method, hence important. The advantage of using dovetail groove is that the O-ring will stay inside during the assembly without falling off.  

 

https://www.youtube.com/watch?v=KJje7GgFBbQ

 

I asked Nor-Cal about in-situ baking to reduce outgassing and they recommended that with viton O-ring seals 200C is max and 150C or lower is safe. I will go with 100 C. Steve doesn’t agree and doesn’t like the idea of baking it during pump down. As per him, once a turbo pump was destroyed (I guess at 40m) after some parts of the O-ring got damaged and got sucked inside. However, I think keeping things at 100 C should be safe as other folks at the cryolab have done the same, successfully.

 

 

 

 

 

 

Attachment 1: bottom_plate_on_spacer.jpg
bottom_plate_on_spacer.jpg
Attachment 2: Crane_dovetail.jpg
Crane_dovetail.jpg
  2264   Wed Nov 28 17:14:11 2018 awadeSummaryTMTFNotes on optimal length mismatch for fiber delay line MZ

I just came across some notes I made about optimal fiber length for the TMTF fiber stabilization scheme.  I'm putting them here for future reference. The idea is to construct a fiber Mach-Zehnder interferometer with one arm much long than the other.  The arm length mismatch gives us a frequency discriminator to which we can lock our laser to.  The question is what length is best, given expected losses, for maximum signal-to-noise?

Choosing the right arm length mismatch

Larger path length mismatch always increases the slope of the ideal MZ interferometer in frequency space.  However, at some point the propagation losses in real fiber reduce the fringe visibility, degrading the signal slope and, therefore, signal-to-noise.

Vinny made some initial calculations that indicated that optimal path length mismatch for SM2000 fiber MZ interferometer was 117 m and in other estimates on order of kilometers. I redid these calculations myself to see if there was an nice analytic expression for the long arm length given a known loss.  It turns out there is:

\Delta L = \frac{1}{\alpha}

Where alpha is the loss per unit length that goes as \textrm{Loss} = \exp(-\alpha L).  For SM2000 fiber, which is the standard type that Thorlabs ship, the losses are estimated at 37.5 dB/km @ 2000 nm.  This makes an alpha of 8.6e-3/m and the estimated optimal fiber mismatch length of 116 m. This agrees with the last number Vinny publish on the elog (ATF:2207).  

I've attached a Mathematica notebook (as a pdf and .nb file) so that we have a common point of reference for future questions of design.  This way we don't have to redo it every time. In the book there is an estimate of the expected frequency equivalent shot noise if we operate the 2 µm MZ with 100 µW of power.  In there I find that, at the shot noise limit, we are looking at order 50 mHz/rtHz noise floor.  This is an order of magnitude worse that the next best standard 1064 nm fiber.  

These calculations were only considering shot noise and not fiber acoustic noise or PD dark noise.  If we up the input power to 1 mW then we should expect a frequency equivalent shot noise of 15 mHz/rtHz.  

Detection at the output

During construction of this experiment for the SURF project Vinny built a single TransImpedance Amplifier (TIA) for each MZ: in loop and out of loop.  We then subtracted a DC offset with a stable voltage reference to place ourselves roughly at the half fringe.  This was ok as a quick dirty first pass. However, the ANU people pointed me to a better way.  We make a subtraction measurement between the two outputs of the MZ.  This doubles the signal slope and to first order is immune to intensity fluctuations of the laser. 

We had trouble locking because when we feedback to the laser current to control frequency the power is also affected.  By making a ballanced measurement this way would make us immune to power changes in the laser and always keep us at the half fringe.  

At this stage the experiment is turned off and unmanned.  When we get another undergraduate we should get them onto make a pair of matched low noise TIA to make a ballanced readout scheme from the MZ.  We need to also look into getting some more PD to be able to also do the out of loop detection at the same time.

Attachment 1: 2umMissmatchWorkings_v1.pdf
2umMissmatchWorkings_v1.pdf 2umMissmatchWorkings_v1.pdf 2umMissmatchWorkings_v1.pdf 2umMissmatchWorkings_v1.pdf 2umMissmatchWorkings_v1.pdf 2umMissmatchWorkings_v1.pdf
Attachment 2: 2umMissmatchWorkings_v1.nb.zip
  2262   Wed Nov 28 12:14:35 2018 awadeDailyProgressWOPOMode field diameter choosing best patch for WOPO input

There is some loss in fiber-to-fiber connection when there is a different mode field diameter between the two.  Ideally one would use the same type of fiber when patching but close enough ussually has lowish losses.  Just to check the numbers, to make sure I'm not getting larger than expected loss going into the WOPO device I looked at the missmatch loss between the single mode patch fibers have have (both single mode polarization maintaining and non-polarization maintaining) and the WOPO's Coastalcon 480PM (1628-10-19) input fiber on the 532 nm side.

For reference

Mismatch in Mode Field Diameter (MFD) between single mode fibers leads to losses at their interface according to

Loss[dB] = 10 log_{10}\left[\frac{4}{\left(\frac{MFD_1}{MFD_2}+\frac{MFD_2}{MFD_1}\right)^2}\right]

 

Some patch fiber parameters

Summary: 488PM-FC-2 (Blue Panda 2 m patch fiber)
Parameter Value
Fiber used PM460-HP
Mode Field Diameter (MFD) 3.3±0.5 @ 515 nm
Propagation loses ≤100 dB/km@ 488 nm
Cutoff wavelength 410±40 nm
Numerical aperture (NA) 0.12
Core (cladding) diameter  3.0 (125 ± 2) µm
Beatlength (of polarization) 1.3 mm @ 460 nm

 

Summary: P3-460B-FC-2  (Yellow SM 2 m patch fiber)
Parameter Value
Fiber used SM450
Mode Field Diameter (MFD) 3.45±0.65 @ 515 nm
Propagation loses ≤50 dB/km@ 488 nm
Cutoff wavelength 410±60 nm
Numerical aperture (NA) 0.12±0.02
Core (cladding) diameter  Not listed (125 ± 2) µm
Beatlength (of polarization) N/A

 

And the fiber input to the AdvR waveguided device

Summary: Coastalcon 480PM (1628-10-19)
Parameter Value
Fiber used Coastalcon 480PM (1628-10-19)
Mode Field Diameter (MFD) 4.0±? @ 515 nm
Propagation loses ≤30 dB/km@ 488 nm
Cutoff wavelength 435±35 nm
Numerical aperture (NA) 0.075±?
Core (cladding) diameter  Not listed (125 ± 2) µm
Beatlength (of polarization) ? Not listed
Extinction Ratio (pol) (dB) ≥25 dB
Insertion loss ≤1.15 dB

 

Mode Field Diameter Mismatch Loss (Between Fibers)

Based on the above specs of mode field diameter the loss due to mismatch is order 0.15 dB for the blue PM460-HP polarization maintaining fiber and -0.1 dB for the singla mode yellow SM450 patch cable.  We can conclude that the mode field dameter mismatch is probably not going to be the dominant loss at the fiber interface with insertion loss likely to be larger at about ~ 1.0dB loss at the plugged interface of the cables.

  2261   Wed Nov 14 16:49:47 2018 RahulUpdateCryo vacuum chamberassembly

The top (seen with several threaded holes along with 16 through holes) and bottom (only 2 threaded holes for lifting and 16 through holes) plate for the vacuum chamber has arrived and I have moved them into the QIL optical bench/table. Using some aluminum struts/bosch, I am making a simple 3’’ tall spacer on top of which the bottom plater will be resting. After wiping them with solvents, I will start assembling the chamber and the plates.

Attachment 1: Bottom_plate.jpg
Bottom_plate.jpg
Attachment 2: top_plate.jpg
top_plate.jpg
  2260   Fri Nov 9 14:57:54 2018 RahulUpdate Cryo-vacuum chamber

SHI Cryocooler – Vacuum chamber assembly

The attached picture shows a schematic of the assembly/connection of SHI cryocooler to the vacuum chamber. The cryocooler has warm flange mounting holes. Using a mating flange – hose/bellows will be connected to the cryocooler. The mating flange will have a port for roughing pump and vacuum gauge connection. The hose/bellows will be connected to the flange reducer which will be bolted to the CF (4-5/8 size) flange of the cryostat.

I will upload a CAD model of the mating hose and will also look for an appropriate size hose and flange reducer.

Attachment 1: Picture1.jpg
Picture1.jpg
  2259   Fri Nov 2 08:59:26 2018 RahulUpdate Vacuum chamber

We have our shiny new vacuum chamber (fabricated by Nor-Cal), now sitting on the optical bench.

Attachment 1: IMG_0005.jpg
IMG_0005.jpg
Attachment 2: IMG_0008.jpg
IMG_0008.jpg
Attachment 3: IMG_0007.jpg
IMG_0007.jpg
  2258   Tue Oct 30 10:17:44 2018 RahulUpdate SHI Cryocooler

We have received the SHI cryocooler CH-104 (figure attached), which has been moved to the QIL. I have inspected all the components after unboxing it. Cold head test report supplied by SHI is attached below.

The cryocooler comes with a HC-4A Zephyr air cooled helium compressor. This compressor is a single stage, air cooled and designed to deliver high pressure helium gas to the cryocooler.

There are 2 helium supply/return (although both the hose says supply, which I am not sure why, hence will check it out) hose along with a kit to install it. This is currently charged with helium pressure of 280 psi, however, once it is installed then the helium pressure has to be adjusted (I am currently reading the manual to assemble the system).

The cryoocoler cold head will be finally placed on a bench which we have bought. I plan to use a breadboard to clamp it down. The compressor will be placed a few feet (based on hose length) away from it. Typically the compressors are noisy, hence later on we can get longer hose to keep the compressors further away.

The vacuum chamber (collar) will be moved in the lab and on the optical bench this Thursday (although this was supposed to be moved in last week, however due inadequate communication by the Caltech moving service this couldn't happen).

The top and bottom flange covers for the vacuum chamber (fabricated by Kurt Lesker) has been shipped on 24 Oct and we should be receiving it this week.

Attachment 1: cryocooler.jpg
cryocooler.jpg
Attachment 2: compressor.jpg
compressor.jpg
Attachment 3: supply_hose.jpg
supply_hose.jpg
Attachment 4: Cold_test_report_SHI.jpg
Cold_test_report_SHI.jpg
  2257   Mon Oct 29 19:05:58 2018 awadeDailyProgressWOPOFixing MM into 532nm collimator

After a number of changes to the path for the pump light (532 nm) it has be very difficult to get any decent amount of pumping power coupled into the waveguide.  I'm re-mode matching light back into the fiber.

The main motivation for the changes was to get a co-propagating 1064 nm beam going into the WOPO so that the non-linear gain could be directly optimized with some seed light.  I installed a dichroic mirror through which the 532 nm is coupled into the fiber and from which 1064 nm is reflected into the coupler from a second path (see picture below).

The fiber is P3-460B-FC-2, but it was just possible to get a couple of µW of power out at 1064 nm out the other end (we probably don't need much).  But after the dichroic was installed and pump path optics were moved around a little to accommodate the new configuration, I found with best pointing the pump light was struggling to get much more than 10% coupling efficiency.

Re mode matching:

I double checked 532 nm beam coming out of the laser with the beam profiler. I confirmed that there was an x-axis waist of 89.0+/-1.2 um at z=-320+/-8 mm and a y-axis fitted waist of 82.7+/-1.0 um at z=-328+/-8 mm (z = 0 is marked on the table after the faraday isolator, the front of the laser head is z = -0.381 m).

Best MM solution was one lens f=350mm@0.37m and a second f=175mm@1.225m.  This was for the F240APC-532 (532 nm) collimator (assuming a 1.48 mm beam diameter).  This is assuming a 460HP fiber which has a slightly different mode size in the fiber; this changes the MM solution slightly but should be close enough.  After a bit of walking of the lenses I found I could bring the 532 nm coupling back up to about 50%.  This can be improved by more walking.

  2256   Mon Oct 22 09:46:10 2018 RahulLab Infrastructure Cryo vacuum chamber setup

The optical bench at the QIL has been re-arranged to accommodate the new cryo vacuum chamber. Today the vacuum chamber (collar fabricated by Nor Cal) will be moved inside the lab, on the optical bench. We have also received the cryocooler from SHI, which will be unboxed and moved inside the lab - to be kept on a separate table. The pumping station (along with the vacuum gauges) has arrived and is currently sitting in the lab. I have filled up the roughing pump with oil and attached a T section (covered up blank flange) to the flange of the tourbopump for a quick start up test. I used a small hose for the exhaust. However, I am in the process of getting a longer hose (12 feet) for a more permanent setup. 

Lab crane has also arrived, which has been assembled (I will attach a pic later on) in the lab.

Attachment 1: lab.jpeg
lab.jpeg
Attachment 2: Guages_and_controller.jpg
Guages_and_controller.jpg
Attachment 3: Crane.jpg
Crane.jpg
Attachment 4: pumping_station.jpg
pumping_station.jpg
  2255   Fri Oct 12 19:05:51 2018 awadeDailyProgressWOPOGluing 532 nm PZT mirror

I'm making a PZT mounted 532 nm mirror, so I have some independant control of the pumping light phase for optimizing the classical gain of the WOPO.

Previously I machined a simple reaction mass out of an old 1 " post (see ATF:2245).  This fits neatly into a standard 1" adjustable mount. To this I have glued a PA44LEW PZT chip and am curing it overnight at 80 C. To this I will stick a broadband visable mirror (Thorlabs BB03-E02) in the next few days and will replace it into the path that I am currently re-modematching into fiber.

Epoxy being used is EPO-TEK 353ND.

ELOG V3.1.3-